INTEGRATING FACTOR


Test for exactness and find the general solution:

1. y(x2 + y)dx + x(x2 − 2y)dy = 0

2. (2y4 + 1)dx + 4xy3dy = 0

3. y(x3y − 1)dx + x(xy4 − 1)dy = 0

4. y2dx + xydy = 0

5. y(y − 1)dx + xdy = 0


1
Expert's answer
2021-06-21T14:51:44-0400

1. "y(x^2 + y)dx + x(x^2 \u2212 2y)dy = 0"


"(x^2y + y^2)dx + (x^3 \u2212 2xy)dy = 0"


"M(x,y)dx+N(x,y)dy=0"


"M(x,y)=x^2y+y^2\implies{\partial M(x,y)\over \partial y}=M_y =x^2 +2y"


"N(x,y)=x^3-2xy\implies{\partial N(x,y)\over \partial x}=N_y =3x^2 -2y"


hence not exact since "M_y \ne N_x"


Integrating Factor = "e^{ \int^x {M_y -N_x\over N}dx }=e^{ \int^x { x^2+2y-3x^2+2y\over x^3-2xy}dx }"


="e^{ \int^x { -2(2y-x^2)\over x(2y-x^2)}dx}=e^{\int -{2\over x}dx}"


="e^{-2\int {1\over x}dx}=e^{lnx^{-2}}={1\over x^2}"


multiplying the differential equation with the integrating factor "{1\over x^2}" gives


"{1\over x^2}[(x^2y + y^2)dx + (x^3 − 2xy)dy = 0]"


"[{x^2y+y^2\over x^2}]dx+[{x^3-2xy\over x^2}]dy=0"


"M(x,y)=y+{y^2\over x^2}\implies{\partial M(x,y)\over \partial y}=M_y =1+{2y\over x^2}"


"N(x,y)=x-{2y\over x}\implies{\partial N(x,y)\over \partial x}=N_y =1 +{2y\over x^2}"


hence exact since "M_y = N_x=1+{2y\over x^2}"


There exist a function "\mu(x,y)" such that


"{\partial \mu(x,y)\over \partial x}=M(x,y)=y+{y^2\over x^2}" and


"{\partial \mu(x,y)\over \partial y}=N(x,y)=x-{2y\over x}"


let "\mu(x,y)=\int^xM(x,y)dx+g(y)"

"=\int^x(y+{y^2\over x^2})dx+g(y)"


"\mu(x,y)=xy-{y^2\over x}+g(y)"


"\implies {\partial \mu(x,y)\over \partial y}=x-{2y\over x}+{\partial g(y)\over \partial y}"


"\implies x-{2y\over x}+{\partial g(y)\over \partial y}=x-{2y\over x}"


"\implies \int g (y)=\int 0dy \implies g(y)=0"


"\implies \mu(x,y)=xy-{y^2\over x}+0"


"\therefore" The general solution is thus "xy-{y^2\over x}=c"


2. "(2y^4 + 1)dx + 4xy^3dy = 0"


"M(x,y)=2y^4+1\implies{\partial M(x,y)\over \partial y}=M_y =8y^3"


"N(x,y)=4xy^3\implies{\partial N(x,y)\over \partial x}=N_y =4y^3"


hence not exact since "M_y \ne N_x"


Integrating Factor = "e^{ \int^x {M_y -N_x\over N}dx }=e^{ \int^x { 8y^3-4y^3\over 4xy^3}dx }=e^{ \int^x { 1\over x}dx }=e^{ lnx}=x"


multiplying the differential equation with the integrating factor "x" gives

"x[(2y^4 + 1)dx + 4xy^3dy = 0]"


"(2xy^4 + x)dx + 4x^2y^3dy = 0"


"M(x,y)=2xy^4+x\implies{\partial M(x,y)\over \partial y}=M_y =8xy^3"


"N(x,y)=4x^2y^3\implies{\partial N(x,y)\over \partial y}=N_y =8xy^3"



hence exact since "M_y = N_x=8xy^2"


There exist a function "\mu(x,y)" such that


"{\partial \mu(x,y)\over \partial x}=M(x,y)=2xy^4+x" and


"{\partial \mu(x,y)\over \partial y}=N(x,y)=4x^2y^3"


let "\mu(x,y)=\int^yN(x,y)dy+g(x)"


"=\int^y4x^2y^3dy+g(x)"


"\mu(x,y)=x^2y^4+g(x)"


"\implies {\partial \mu(x,y)\over \partial x}=2xy^4+{\partial g(x)\over \partial x}"


"\implies 2xy^4+{\partial g(x)\over \partial x}=2xy^4+x"


"\implies \int d g (x)=\int xdx \implies g(x)={x^2\over 2}"


"\implies \mu(x,y)=x^2y^4+{x^2\over 2}"


"\therefore" The general solution is thus "x^2y^4+{x^2\over 2}=c"

3."y(x^3y \u2212 1)dx + x(xy^4 \u2212 1)dy = 0"


"(x^3y ^2\u2212 y)dx + (x^2y^4 \u2212 x)dy = 0"


"M(x,y)=x^3y ^2− y\implies{\partial M(x,y)\over \partial y}=M_y =2x^3y-1"


"N(x,y)=x^2y^4 − x\implies{\partial N(x,y)\over \partial x}=N_x =2xy^4-1"


hence not exact since "M_y \ne N_x"


The integrating factor of the form p(x) or p(y) do not exist for the differential equation hence has no solution


4."y^2dx + xydy = 0"


"M(x,y)=y^2\implies{\partial M(x,y)\over \partial y}=M_y =2y"


"N(x,y)=xy\implies{\partial N(x,y)\over \partial x}=N_x =y"


hence not exact since "M_y \ne N_x"


Integrating Factor = "e^{ \int^x {M_y -N_x\over N}dx }=e^{ \int^x { 2y-y\over xy}dx }=e^{ \int^x { 1\over x}dx }=e^{ lnx}=x"


multiplying the differential equation with the integrating factor "x" gives


"x[ y^2dx + xydy = 0]"


"xy^2dx + x^2ydy = 0"


"M(x,y)=xy^2\implies{\partial M(x,y)\over \partial y}=M_y =2xy"


"N(x,y)=x^2y\implies{\partial N(x,y)\over \partial x}=N_x =2xy"


hence exact since "M_y = N_x=2xy"


There exist a function "\mu(x,y)" such that


"{\partial \mu(x,y)\over \partial x}=M(x,y)=xy^2" and


"{\partial \mu(x,y)\over \partial y}=N(x,y)=x^2y"


let "\mu(x,y)=\int^xM(x,y)dx+g(y)"


"=\int^xxy^2dy+g(x)"


"\mu(x,y)={x^2\over 2}y^2+g(x)"


"\implies {\partial \mu(x,y)\over \partial y}=x^2y+{\partial g(y)\over \partial y}"


"\implies x^2y+{\partial g(y)\over \partial y}=x^2y"


"\implies \int d g (y)=\int 0dx \implies g(x)=0"


"\implies \mu(x,y)={1\over 2}x^2y^2+0"


"\therefore" The general solution is thus "{1\over 2}x^2y^2=c"


5. "y(y \u2212 1)dx + xdy = 0"


"(y^2 \u2212 y)dx + xdy = 0"


"M(x,y)=y^2-y\implies{\partial M(x,y)\over \partial y}=M_y =2y-1"


"N(x,y)=x\implies{\partial N(x,y)\over \partial x}=N_x =1"


hence not exact since "M_y \ne N_x"


Integrating Factor = "e^{ \int^y {M_y -N_x\over N}dy }=e^{ \int^y { 1-2y+1\over y^2-y}dy }=e^{ \int^y{ 2-2y\over y^2-y}dy }"


evaluate "{ \int^y{ 2-2y\over y^2-y}dy }" by partial fraction


"{ 2-2y\over y(y-1)}={A\over y}+{B\over y-1}"


"{ 2-2y\over y(y-1)}={A(y-1)+By\over y(y-1)}"


"\implies 2-2y=Ay-A+By=(A+B)y-A"


equating coefficients of y terms together and constants together we have


"A+B=-2 and A=-2\implies B=0"


"\therefore{ 2-2y\over y(y-1)}={-2\over y}+{0\over y-1}={-2\over y}"


"\therefore { \int^y{ 2-2y\over y^2-y}dy }=-2\int{1\over y}dy=-2lny=ln y^{-2}"


"\therefore" Integrating Factor "=e^{lny^{-2}}={1\over y^2}"


multiplying the differential equation with the integrating factor "{1\over y^2}" gives


"{1\over y^2}[ (y^2 − y)dx + xdy = 0]"


"{y-1\over y}dx + {x\over y^2}dy = 0"


"M(x,y)={y-1\over y}\implies{\partial M(x,y)\over \partial y}=M_y ={1\over y^2}"


"N(x,y)={x\over y^2}\implies{\partial N(x,y)\over \partial x}=N_x ={1\over y^2}"


hence exact since "M_y = N_x={1\over y^2}"


There exist a function "\mu(x,y)" such that


"{\partial \mu(x,y)\over \partial x}=M(x,y)={y-1\over y}" and


"{\partial \mu(x,y)\over \partial y}=N(x,y)={x\over y^2}"


let "\mu(x,y)=\int^xM(x,y)dx+g(y)"


"=\int^x{y-1\over y}dx+g(y)"


"\mu(x,y)=x{y-1\over y}+g(y)"


"\implies {\partial \mu(x,y)\over \partial y}={x\over y^2}+{\partial g(y)\over \partial y}"


"\implies {x\over y^2}+{\partial g(y)\over \partial y}={x\over y^2}"


"\implies \int d g (y)=\int 0dy \implies g(y)=0"


"\implies \mu(x,y)=x{y-1\over y}+0"


"\therefore" The general solution is thus "x{y-1\over y}=c"